# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v4

@@ Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, one universal threshold could fail to account for distributional variations by subgroups. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts. These discrepancies can lead to misclassifications that disproportionately affect certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors. We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy. In experiments with nine AI text classifiers on three datasets, FairOPT decreases overall balanced error rate (BER) discrepancy by 12% while minimally sacrificing accuracy by 0.003%. Our framework paves the way for more robust classification in AI-generated content detection via post-processing. We release our data, code, and project information at URL.
1 Introduction
@@
Figure 2: Applying adaptive thresholds to different probability distributions by subgroup in the test dataset, as seen through cumulative density functions (CDFs). We partition texts by three length categories and five personality traits as noted in Table 1. Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold using AUROC-based method (in gray) do not account for subgroup-specific distributional variations. We used fifteen thresholds that have been generated using FairOPT to address the varying probabilistic distribution from AI text classifiers. The training is based on the RoBERTa-large detector. Details about the thresholds are in Appendix E.
@@
To overcome these limitations, we introduce FairOPT, a group-adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1). Our method balances performance (e.g., accuracy (ACC), F1, precision) and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds. Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces BER disparity across subgroups while incurring only a marginal tradeoff in overall performance.
To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving a 12% reduction in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
2 Related Work
AI text detectors:
AI text detectors have been developed to tackle issues related to misinformation, integrity, and security. RADAR utilizes adversarial learning techniques [13], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e. the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [14], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [15], and RoBERTa-based detectors fine-tune pretrained language models [16]. We employed probabilistic classifiers for an AI-likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
Static thresholding for classification:
The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [8]. This default rule is commonly used in many fields like photogrammetry [17], ecology [18, 19], and machine learning [20]. Nevertheless, as highlighted by Freeman and Moisen [8], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
Universal thresholding with optimization:
@@ 2 Related Work
Adaptive thresholding across groups:
Several studies advocate for group-adaptive decision thresholds to mitigate disparities [12, 27]. Menon and Williamson [28] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy. Similarly, Corbett-Davies et al. [29] demonstrate that group-specific thresholds can facilitate fairness-aware classification. Canetti et al. [30] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. Jang et al. [12] present threshold-adaptation methods to ensure fair classification, and Bakker et al. [27] explore threshold tuning to maintain stable classification performance. Building on these foundations, our framework FairOPT introduces a novel group-adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
@@
3 Conceptual Framework and Methodology
@@ 3.1 Objective
3.1.1 Preliminaries of AI Text Detection
The detection of AI-generated text is commonly formulated as a probabilistic classification for a binary classification, where the input space $\mathcal{X}$ consists of the sequence of tokens derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written and $1$ for AI-generated. To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated [16, 13]. This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$. Thus, the predicted label, AI-written, $\hat{Y}\in\{0,1\}$ is determined as follows:
@@ 3.1.1 Preliminaries of AI Text Detection
3.1.2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix B.1.
3.1.3 Fairness metrics
To quantify disparity, we use Demographic Parity (DP) [31] and Balanced Error Rate (BER) [32]. DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equal across all groups. We define the DP metric:
@@ 3.1.3 Fairness metrics
In our subgroup-adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP. For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
@@
Details about fairness metrics are noted in Appendix B.2.
Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section 3.2. Details about the tradeoff are in Appendix H.2.
3.2 Relaxed Fairness
@@
[…]
@@
3.3 Group-Adaptive Threshold Optimization
@@
where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
@@
[…]
@@
By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria. For full implementation details and hyperparameter specifications, refer to Appendix H.1.
@@
Figure 6: Threshold optimization based on the AUROC-based method under an $FPR\leq 10\%$ constraint (see Appendix C). The ROC curve (blue line) illustrates the trade-off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
4 Experiments
4.1 Dataset
Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [36], MAGE [37], and SemEval24 [38] with an even balance between human and AI text to avoid class imbalance. The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles. This experiment employed 19,117 instances for the training and 4,779 samples for the testing. We conducted the featured engineering noted in 4.3 and AI detections noted in 4.2 to extract features and get inferences from classifiers.
4.2 AI-Generated Text Classifiers
To classify AI-generated text, we employed multiple AI text detectors, DeskLib [39], E5LoRA [40], FastDetectGPT [41], RADAR [13], RoBERTa-base and -large [16], and SimpleAI’s detector [42] for open-source availability. Among closed-source models, GPT4.1-nano and GPT4o-mini [43, 44] have been used for LLM-based text detection [7]. These detectors were selected based on their common application and used in both the train and test datasets. The closed-source text detection is only conducted in the test dataset.
Figure 7: FairOPT decreases the Balanced Error Rate (BER) discrepancy of AI detectors compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$). The error bar shows the standard deviation from the mean.
4.3 Subgroup Attributes
We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics. These attributes are shown in Table 1. We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500). We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [45]. Details are noted in Appendix F. These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
@@ 4.3 Subgroup Attributes
Table 1: Subgroup attributes from the feature engineering.
@@
Figure 8: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4. Our method incurs minimal loss of ACC, achieving $\mu=0.6344\pm 0.1518$ compared to Static ($\mu=0.6376\pm 0.1975$) and ROCFPR ($\mu=0.6249\pm 0.2057$) methods.
4.4 Discovery of Subgroup Bias in Text Detection

### v5

# Minseok Jung1, Cynthia Fuertes Panizo2, Liam Dugan3, Yi R. (May) Fung1, Pin-Yu Chen4, Paul Pu Liang1

@@ Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, one universal threshold could fail to account for distributional variations by subgroups. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts. These discrepancies can lead to misclassifications that disproportionately affect certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors. We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy. FairOPT showed notable discrepancy mitigation across nine detectors and three heterogeneous datasets, and the remarkable mitigation of the minimax problem by decreasing overall discrepancy 27.4% across five metrics while minimally sacrificing accuracy by 0.005%. Our framework paves the way for more robust classification in AI-generated content detection via post-processing. We release our data, code, and project information at URL.
I Introduction
@@
To overcome these limitations, we introduce FairOPT, a group‐adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1). Our method balances performance and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds. Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces discrepancy across subgroups while incurring only a marginal tradeoff in overall performance.
To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving overall 27.4% reduction across five metrics in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
II Literature review
AI text detectors
AI text detectors have been developed to tackle issues related to misinformation, integrity, and security. RADAR utilizes adversarial learning techniques [24], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e. the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [18], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [35], and RoBERTa-based detectors fine-tune pretrained language models [42]. We employed probabilistic classifiers for an AI‐likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
Static thresholding for classification
The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [17]. This default rule is commonly used in many fields like photogrammetry [41], ecology [32, 21], and machine learning [31]. Nevertheless, as highlighted by [17], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
Universal thresholding with optimization
@@ II Literature review
Adaptive thresholding across groups
Several studies advocate for group‐adaptive decision thresholds to mitigate disparities [25, 2]. [33] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy. Similarly, [9] demonstrate that group-specific thresholds can facilitate fairness-aware classification. [7] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. [25] present threshold-adaptation methods to ensure fair classification, and [2] explore threshold tuning to maintain stable classification performance. Building on these foundations, our framework FairOPT introduces a novel group‐adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
III Conceptual Framework and Methodology
Figure 2: FairOPT decreases discrepancy of AI detectors compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.613$, outperforming ROCFPR ($\mu=0.920$) and Static ($\mu=0.887$). The error bar shows the standard deviation from the mean. Each bar shows average of five types of discrepancies that has been introduced in III-A4.
@@ III Conceptual Framework and Methodology
TABLE I: FairOPTsignificantly reduces subgroup discrepancies across all evaluation metrics while keeping accuracy on par with baselines. This table compares thresholding methods: FairOPT, ROCFPR, and Static on ACC and five fairness metrics that measure discrepancies.
Metric FairOPT ROCFPR Static
ACC 0.6341 0.6249 0.6376
MaxBER Discrepancy 0.5852 0.7592 0.7507
Demographic Parity Difference 0.6772 0.9826 0.9730
Equalized Odds Difference 0.6379 0.9198 0.8501
FPR Discrepancy 0.4313 0.9715 0.9471
Predictive Parity Difference 0.7333 0.9684 0.9130
TABLE II: FairOPT achieves the lowest subgroup discrepancies among three thresholding methods while maintaining accuracy. Comparison of thresholding methods has been made across three datasets: MAGE (MG), RAID (RD), and SemEval (SE) and their overall averages (All). MaxBER = BER Discrepancy, MaxDP = Demographic Parity Discrepancy, MaxEO = Equalized Odds Discrepancy, MaxFPR = False Positive Rate Discrepancy, MaxPP = Predictive Parity Discrepancy. The discrepancy based on the MiniMax problem has been noted in III-A4.
Method MaxBER Disc. MaxDP Disc. MaxEO Disc.
All MG RD SE All MG RD SE All MG RD SE
Static 0.751 0.730 0.771 0.754 0.973 0.972 0.958 0.987 0.850 0.859 0.838 0.852
ROCFPR 0.759 0.745 0.783 0.753 0.983 0.978 0.977 0.992 0.920 0.949 0.901 0.906
FairOPT 0.585 0.582 0.582 0.591 0.677 0.671 0.659 0.700 0.638 0.635 0.629 0.649
Method MaxFPR Disc. MaxPP Disc. Accuracy
All MG RD SE All MG RD SE All MG RD SE
Static 0.947 0.939 0.974 0.932 0.913 0.902 0.920 0.918 0.638 0.639 0.636 0.638
ROCFPR 0.972 0.971 0.985 0.961 0.968 0.967 0.982 0.958 0.625 0.624 0.625 0.626
FairOPT 0.431 0.444 0.416 0.432 0.733 0.746 0.713 0.738 0.634 0.633 0.628 0.640
@@ III-A Objective
III-A1 Preliminaries of AI Text Detection
[…]
@@ III-A1 Preliminaries of AI Text Detection
III-A2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix -A.
III-A3 Minimax problem
We quantify subgroup disparity using five complementary metrics—Demographic Parity (DP) [26], Balanced Error Rate (BER) [16], Equalized Odds, FPR, and Predictive Parity. For each metric $k$, let $m_{k,g}$ denote the value on subgroup $g$; we define the inter-group discrepancy $\Delta_{k}\!=\!\max_{g}m_{k,g}\;-\;\min_{g}m_{k,g}$ (lower is better). We then adopt a minimax objective,
$\min_{\theta}\;\max_{k}\;\Delta_{k}(\theta),$
because fairness guarantees should hold in the worst subgroup rather than only on average. Controlling $\max_{k}\Delta_{k}$ addresses tail risk and improves robustness.
III-A4 Fairness metrics
DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equal across all groups. We define the DP metric:
@@ III-A4 Fairness metrics
In our subgroup‐adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP. For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
@@
Discrepancy has been measured based on minimax problem to minimize the maximum subgroup discrepancy (e.g., BER gap, FPR gap), thereby reducing the worst-case disparity.
Equalized Odds (EO) requires that a classifier’s predictions are conditionally independent of sensitive attributes $S$ given the true label $Y$ [22]. This means that both the TPR and FPR should be equal across groups. In practice, EO measures whether individuals from different subgroups are treated similarly in terms of both favorable outcomes (correctly identifying positives) and unfavorable outcomes (incorrectly labeling negatives as positives). The discrepancy is defined as the maximum gap across subgroups in either TPR or FPR:
$\Delta_{\text{EO}}$ $=\max\Bigl(\max_{g}\text{TPR}-\min_{g}\text{TPR},\;\max_{g}\text{FPR}-\min_{g}\text{FPR}\Bigr).$ (5)
Perfect fairness under EO is achieved when $\Delta_{\text{EO}}=0$, meaning that subgroup differences in both detection sensitivity and error rates are eliminated.
False Positive Rate (FPR) Discrepancy focuses specifically on subgroup differences in the rate of false alarms, i.e., when the classifier predicts $\hat{Y}=1$ despite the true label being $Y=0$. The discrepancy is formally defined as the difference between the maximum and minimum false positive rates across groups:
$\Delta_{\text{FPR}}$ $=\max_{g}\text{FPR}_{g}-\min_{g}\text{FPR}_{g}.$ (6)
This metric captures whether some groups are disproportionately subjected to incorrect positive predictions, which is critical in high-stakes applications such as fraud detection.
Predictive Parity (PP) Discrepancy measures whether a positive prediction carries the same meaning for different groups, i.e., whether $P(Y=1\mid\hat{Y}=1,S=g)$ is group-independent. The discrepancy is defined as:
$\Delta_{\text{PP}}$ $=\max_{g}\text{PP}_{g}-\min_{g}\text{PP}_{g},$ (7)
where $\text{PP}_{g}=\frac{\text{TP}_{g}}{\text{TP}_{g}+\text{FP}_{g}}$. PP evaluates whether the precision of a classifier, defined as the proportion of true positives among all predicted positives, is consistent across subgroups.
Addressing the mini–max problem in fairness is important because it explicitly targets the worst-case disparity across subgroups rather than optimizing only for average performance. Details about fairness metrics are noted in Appendix -B.
Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section III-B. Details about the tradeoff are in Appendix -H.
III-B Relaxed Fairness
@@
6:   if ($\mathrm{ACC}_{i}\geq\alpha$ and $\mathrm{F1}_{i}\geq\beta$ for all ii) and (fairness gap $\leq\delta_{\text{fair}}$) then
@@
III-C Group-Adaptive Threshold Optimization
@@
where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $\cdot$$_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
@@
If $\Delta_{\theta}$ falls below a tolerance $\epsilon_{\mathrm{tol}}$ and all groups meet the performance criterion ($\mathrm{ACC}_{i}\geq\alpha$, $\mathrm{F1}_{i}\geq\beta$) and fairness disparities ($\Delta_{k}\leq\delta_{\mathrm{fair}}$) hold, the algorithm terminates.
@@
By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria. For full implementation details and hyperparameter specifications, refer to Appendix -G.
@@
Figure 6: Threshold optimization based on the AUROC‐based method under an $FPR\leq 10\%$ constraint (see Appendix AUROC). The ROC curve (blue line) illustrates the trade‐off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
IV Experiment
IV-A Dataset
Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [12], MAGE [28], and SemEval24 [45] with an even balance between human and AI text to avoid class imbalance. The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles. This experiment employed 19,117 instances for the training and 4,779 samples for the testing. We conducted the featured engineering noted in IV-C and AI detections noted in IV-B to extract features and get inferences from classifiers.
IV-B AI-Generated Text Classifiers
To classify AI-generated text, we employed multiple AI text detectors, DeskLib [10], E5LoRA [50], FastDetectGPT [3], RADAR [24], RoBERTa-base and -large [42], and SimpleAI’s detector [19] for open-source availability. Among closed-source models, GPT4.1-nano and GPT4o-mini [38, 37] have been used for LLM-based text detection [5]. These detectors were selected based on their common application and used in both the train and test datasets. They are based on probabilistic inference framework that needs a threshold for the final classification. The closed-source text detection is only conducted in the test dataset.
IV-C Subgroup Attributes
We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics. These attributes are shown in Table III. We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500). We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [34]. The third feature is formality that measures the formal expression of the sentence based on Heylighen & Dewaele’s formality score [23] and two categories (formal and informal) are labeled. The fourth feature is a binary sentiment label that has been inferred using DistilBERT-based SST-2 classifier[40]. Details are noted in Appendix Feature Engineering. These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
TABLE III: Subgroup attributes from the feature engineering.
@@ IV-C Subgroup Attributes
Formality Formality of the writing Formal, informal (k=2)
Sentiment Binary sentiment Positive, negative (k=2)
@@
Figure 7: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4. Our method incurs minimal loss of ACC, achieving $\mu=0.6341$ compared to Static ($\mu=0.6249$) and ROCFPR ($\mu=0.6376$) methods.
IV-D Discovery of Subgroup Bias in Text Detection